OpenStudy (anonymous):

Identify the slope and y-intercept of the line with the given info: 1. y=-9/7x+2 2. 9x+3y=6 Write an equation in slope intercept form of the line that passes through the given point and has the given slope m. 3. (1,3); m=4 4. (-2,5); m=-3

OpenStudy (anonymous):

please helppp :(

OpenStudy (chihiroasleaf):

the general form of equation of line is \[y=mx+c\] with m is the slope of the line and c is the y-intercept of the line for the first number,can you identify what is m and c? for the second, you should write down the equation in the form y = mx+c if a line pass through point \((x_1, y_1)\) and the slope of the line is \(m\) ,then the equation of the line is \[y-y_1 = m(x-x_1)\] you just need to plug in your information into this formula and simplify

OpenStudy (anonymous):

If you do it for me, Ill give u a medal and fan u